# Novel duality-invariant theories of electrodynamics

### v1

@@ Abstract
But one could consider other models of electrodynamics fulfilling Lorentz and gauge invariance beyond Maxwell theory. These are called theories of non-linear electrodynamics (NED) and have been extensively studied for a long time. Some prominent models include Born-Infeld theory [1] — introduced to provide a finite self-energy for the electric field of a charged point particle — or the Euler-Heisenberg Lagrangian [2], incorporating one-loop quantum corrections — for other NED models and additional background on the topic, we refer the reader to, e.g., [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. Nonetheless, if one really takes symmetries as the true guiding principle for the construction of physical theories, nothing prevents us from considering other theories of NED which do not violate other fundamental principles of physics (such as the absence of ghosts or causality).
@@
^$\frac{\partial\hat{\mathcal{L}}_{\gamma}}{\partial s}=\frac{\partial\mathcal{L}_{0}}{\partial z}\frac{\partial z}{\partial s}\,,\quad\frac{\partial\hat{\mathcal{L}}_{\gamma}}{\partial p}=\frac{\partial\mathcal{L}_{0}}{\partial z}\frac{\partial z}{\partial p}+\frac{\partial\mathcal{L}_{0}}{\partial p}\,,$ (10)
@@
one directly learns that ^$\hat{\mathcal{L}}_{\gamma}$ satisfies (6) with the same $\psi$ as $\mathcal{L}_{0}$ and we conclude. ∎
@@
±$\mathcal{L}_{\pm}^{\mathrm{GR},\psi}=\mathfrak{f}_{K}(\sqrt{u}+\sqrt{v},0)\pm\mathfrak{f}_{K}(\sqrt{u}-\sqrt{v},\mathrm{sign}(p)\,\psi),$ (21)
$\mathfrak{f}_{K}(b,c)=\frac{b}{2}\sqrt{2(K^{2}-c)-b^{2}}$
@@
±$\pm\mathfrak{f}_{K}(e^{-\gamma/2}\sqrt{u}-e^{\gamma/2}\sqrt{v},{\rm sign}(p)\,\psi)\,.$
@@
where $\ell(\tau)$ is an arbitrary function of $\tau\geq 0$ and ˙=d/$\dot{\ell}=\mathrm{d}\ell/d\tau$. Indeed, if the Lagrangian may be expressed as (25), then
@@
where $\dot{\ell}$ is evaluated at $\tau(e^{-\gamma}u,e^{\gamma}v)$ and $\dot{q}$ at $\sigma(u,v)$. As a consequence, we conclude that
@@
Note that we can refine a bit more Theorem 2. Indeed, assume a theory $\mathcal{L}$ that is expressed as in (25) and consider that ˙˙$\dot{\ell}\geq\dot{\ell}_{0}$ for a real number ˙$\dot{\ell}_{0}>0$. Then, $\mathcal{M}_{\gamma}(\mathcal{L})$ with $\gamma\geq-\log(\dot{\ell}_{0})$ will be causal. If ˙$\dot{\ell}_{0}>1$, negative values of $\gamma$ are allowed. If ˙$\dot{\ell}_{0}<1$ — what implies that the starting theory was not causal —, the ModMax map allows one to construct causal generalizations of the seed theory for $\gamma\geq-\log(\dot{\ell}_{0})$.
@@
so that the equations for the additional fields conforming $\psi$ will not transform in general covariantly. A possible resolution of this problem would be to ensure either that $\dfrac{\partial\mathcal{L}}{\partial\psi}$ always appears in the equations under no derivation operation or, instead, that this term is always affected by derivatives — in this way, (34) guarantees covariance of the equations. The second case may be achieved if $\psi$ is given by a total derivative. A particular example of this situation takes place if [^6] Interestingly enough, the subsequent duality-invariant theory would also be conformal in this case. $\psi=\mathfrak{p}$, where $-4\mathfrak{p}=\mathcal{F}_{ab}\left(\star\mathcal{F}\right)^{ab}$ is the invariant associated with the field strength $\mathcal{F}_{ab}=2\partial_{[a}\mathcal{A}_{b]}$ of a different gauge field $\mathcal{A}_{a}$.
Beyond the possible interpretations for $\psi$, let us now focus on the study of solutions of (6) with non-trivial $\psi$. Its direct resolution conforms a daunting task, so it is convenient to look for special solutions of (6) which satisfy additional algebraic and differential properties. Specifically, we will restrict to theories $\mathcal{L}=\mathcal{L}(s,p,\psi)$ which are homogeneous functions of degree one, i.e., $\mathcal{L}(\Theta s,\Theta p,\Theta\psi)=\Theta\mathcal{L}$ for any function $\Theta$, so that
@@
$-p\left(\frac{\partial\mathcal{L}}{\partial p}\right)^{2}+\psi=p\,,$ (36)
@@
Define $y=\dfrac{p}{\psi}$. In the weak-field limit $01>>y>$ 0 and disregarding the term $\mathfrak{h}(\psi)$, $\mathcal{L}_{\rm gBB}$ reduces to
@@
$\mathcal{L}_{\psi{\rm M}}(s,p,\psi)=\frac{s}{\sqrt{2}}\mathfrak{t}_{+}+\frac{p}{\sqrt{2}}\mathfrak{t}_{-}-\frac{\psi}{2}\mathfrak{t}_{\psi}\,,$ (40)
$\mathfrak{t}_{+}=\sqrt{1-\frac{\psi\,p}{s^{2}+p^{2}}+\sqrt{\frac{s^{2}+(\psi-p)^{2}}{s^{2}+p^{2}}}}\,,$ (41)
$\mathfrak{t}_{-}=\frac{1}{\psi s}\sqrt{\psi^{2}s^{2}\left[\frac{\psi p}{s^{2}+p^{2}}-1+\sqrt{\frac{s^{2}+(\psi-p)^{2}}{s^{2}+p^{2}}}\right]}\,,$ (42)
$\mathfrak{t}_{\psi}=\arctan\left(\frac{s}{p}\right)+2\arctan\left(\frac{\mathfrak{t}_{-}}{\sqrt{2}+\mathfrak{t_{+}}}\right)\,.$ (43)
@@
$+\frac{\psi\left[(y+\sqrt{x^{2}+y^{2}})\omega^{+}+|x|\omega^{-}\right]}{\sqrt{2}(x^{2}+y^{2})^{3/4}\,\mathrm{sign}(x)}+\mathcal{O}\left(x^{3/2},y^{3/2}\right)\,,$ (48)
@@ Appendix A Duality-invariance condition with $\psi\neq 0$ in the Hamiltonian formalism
$\dot{\mathbf{B}}$ =×$=-\mathbf{\nabla}\times\mathbf{E}\,,\quad\mathbf{\nabla}\cdot\mathbf{B}=0\,,$ (51)
$\dot{\mathbf{D}}$ =×$=\mathbf{\nabla}\times\mathbf{H}\,,\quad\mathbf{\nabla}\cdot\mathbf{D}=0\,,$ (52)
@@
Appendix B Solving (6) and (35) perturbatively

### v3

@@ Abstract
But one could consider other models of electrodynamics fulfilling Lorentz and gauge invariance beyond Maxwell theory. These are called theories of non-linear electrodynamics (NED) and have been extensively studied for a long time. Some prominent models include Born-Infeld theory [1] — introduced to provide a finite self-energy for the electric field of a charged point particle — or the Euler-Heisenberg Lagrangian [2], incorporating one-loop quantum corrections — for other NED models and additional background on the topic, we refer the reader to, e.g., [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. Nonetheless, if one really takes symmetries as the true guiding principle for the construction of physical theories, nothing prevents us from considering other theories of NED which do not violate other fundamental principles of physics (such as the absence of ghosts or causality).
@@
$\frac{\partial\hat{\mathcal{L}}_{\gamma}}{\partial s}=\frac{\partial\mathcal{L}_{0}}{\partial z}\frac{\partial z}{\partial s}\,,\quad\frac{\partial\hat{\mathcal{L}}_{\gamma}}{\partial p}=\frac{\partial\mathcal{L}_{0}}{\partial z}\frac{\partial z}{\partial p}+\frac{\partial\mathcal{L}_{0}}{\partial p}\,$, (10)
@@
one directly learns that $\hat{\mathcal{L}}_{\gamma}$ satisfies (6) with the same $\psi$ as $\mathcal{L}_{0}$ and we conclude. ∎
@@
$\hskip-5.69046pt\mathcal{L}_{\pm}^{\mathrm{GR},\psi}=\mathfrak{f}_{K}(\sqrt{u}+\sqrt{v},0)\pm\mathfrak{f}_{K}(\sqrt{u}-\sqrt{v},\mathrm{sign}(p)\,\psi),\hskip-2.84544pt$ (21)
$\hskip-5.69046pt\mathfrak{f}_{K}(b,c)=\frac{b}{2}\sqrt{2(K^{2}-c)-b^{2}}$
@@
$\pm\mathfrak{f}_{K}(e^{-\gamma/2}\sqrt{u}-e^{\gamma/2}\sqrt{v},{\rm sign}(p)\,\psi)\,.$
@@
where $\ell(\tau)$ is an arbitrary function of $\tau\geq 0$ and $\dot{\ell}=\mathrm{d}\ell/d\tau$. Indeed, if the Lagrangian may be expressed as (25), then
@@
[…]
@@
Note that we can refine a bit more Theorem 2. Indeed, assume a theory $\mathcal{L}$ that is expressed as in (25) and consider that $\dot{\ell}\geq\dot{\ell}_{0}$ for a real number $\dot{\ell}_{0}>0$. Then, $\mathcal{M}_{\gamma}(\mathcal{L})$ with $\gamma\geq-\log(\dot{\ell}_{0})$ will be causal. If $\dot{\ell}_{0}>1$, negative values of $\gamma$ are allowed. If $\dot{\ell}_{0}<1$ — what implies that the starting theory was not causal —, the ModMax map allows one to construct causal generalizations of the seed theory for $\gamma\geq-\log(\dot{\ell}_{0})$.
@@
so that the equations for the additional fields conforming $\psi$ will not transform in general covariantly. A possible resolution of this problem would be to impose that $\psi$ is given by a total derivative term. In such a case, the relevant dangerous terms $\dfrac{\partial\mathcal{L}}{\partial\psi}$ will always be affected by derivatives in the equations for the fields composing $\psi$. As a result, (34) will guarantee the invariance of the equations. Various possible choices for $\psi$ that correspond to total derivatives and provide equations for the fields composing $\psi$ that are invariant under duality transformations are presented in the Supplemental Material.
Let us now study solutions of (6) with non-trivial $\psi$. Its direct resolution conforms a daunting task, so it is convenient to look for special solutions with additional properties. Particularly, we will restrict to theories $\mathcal{L}=\mathcal{L}(s,p,\psi)$ which are homogeneous functions of degree one, i.e., $\mathcal{L}(\Theta s,\Theta p,\Theta\psi)=\Theta\mathcal{L}$ for any function $\Theta$, so that
@@
$p\left(\frac{\partial\mathcal{L}}{\partial p}\right)^{2}+\psi=p\,$, (36)
@@
Define $y=\dfrac{p}{\psi}$. In the weak-field limit $1>>y>0$ and disregarding the term $\mathfrak{h}(\psi)$, $\mathcal{L}_{\rm gBB}$ reduces to
@@
$\hskip-5.69046pt\mathcal{L}_{\psi{\rm M}}(s,p,\psi)=\frac{s}{\sqrt{2}}\mathfrak{t}_{+}+\frac{p}{\sqrt{2}}\mathfrak{t}_{-}-\frac{\psi}{2}\mathfrak{t}_{\psi}\,$, (40)
$\hskip-5.69046pt\mathfrak{t}_{+}=\sqrt{1-\frac{\psi\,p}{s^{2}+p^{2}}+\sqrt{\frac{s^{2}+(\psi-p)^{2}}{s^{2}+p^{2}}}}\,$, (41)
$\hskip-5.69046pt\mathfrak{t}_{-}=\frac{1}{\psi s}\sqrt{\psi^{2}s^{2}\left[\frac{\psi p}{s^{2}+p^{2}}-1+\sqrt{\frac{s^{2}+(\psi-p)^{2}}{s^{2}+p^{2}}}\right]}\,$, (42)
$\hskip-5.69046pt\mathfrak{t}_{\psi}=\arctan\left(\frac{s}{p}\right)+2\arctan\left(\frac{\mathfrak{t}_{-}}{\sqrt{2}+\mathfrak{t_{+}}}\right)\,.$ (43)
@@
$\hskip-28.45274pt+\frac{\psi\left[(y+\sqrt{x^{2}+y^{2}})\omega^{+}+|x|\omega^{-}\right]}{\sqrt{2}(x^{2}+y^{2})^{3/4}\,\mathrm{sign}(x)}+\mathcal{O}\left(x^{3/2},y^{3/2}\right)\,$, (48)
@@ Appendix A Duality-invariance condition with $\psi\neq 0$ in the Hamiltonian formalism
$\dot{\mathbf{B}}$ $=-\mathbf{\nabla}\times\mathbf{E}\,,\quad\mathbf{\nabla}\cdot\mathbf{B}=0\,$, (51)
$\dot{\mathbf{D}}$ $=\mathbf{\nabla}\times\mathbf{H}\,,\quad\mathbf{\nabla}\cdot\mathbf{D}=0\,$, (52)
@@
Appendix B Instances of dynamical choices for $\psi$
Let us consider a certain theory satisfying the duality-invariance condition (6). As explained in the body of the manuscript, allowing $\psi$ to be constructed from other dynamical fields entails an important danger: the equations of motion for such additional fields might not be invariant under duality transformations. One may solve this problem if $\psi$ corresponds to a total derivative term, as it is illustrated below with three different examples.
  1. 1.
    $\psi=\lambda\Box\Phi$, where $\Phi$ is a scalar field, $\Box$ stands for the Laplacian operator and $\lambda$ a certain coupling with units of length-1. Adding the scalar kinetic term, one could consider the scalar-vector theory:
    $\tilde{\mathcal{L}}_{F_{ab},\Phi}=\frac{\beta}{2}\Phi\Box\Phi+\mathcal{L}(s,p,\lambda\Box\Phi)\,$, (58)
    where $\beta$ is a dimensionless coupling and $\mathcal{L}$ satisfies (6) with $\psi=\lambda\Box\Phi$. The equation of motion for the scalar reads:
    $\beta\Box\Phi+\lambda\Box\left(\frac{\partial\mathcal{L}}{\partial\psi}\right)=0\,.$ (59)
    Since $\frac{\partial\mathcal{L}}{\partial\psi}$ is necessarily affected by derivatives in (59), (34) ensures that (59) remains indeed invariant under electromagnetic duality rotations. Nonetheless, the equations will feature higher derivatives.
  2. 2.
    $\psi=\lambda\mathfrak{p}$, where $\lambda$ is a dimensionless coupling and $\mathfrak{p}=-\frac{1}{4}\mathcal{F}_{ab}\left(\star\mathcal{F}\right)^{ab}$ for a field strength $\mathcal{F}_{ab}=2\partial_{[a}\mathcal{A}_{b]}$ of a different gauge field $\mathcal{A}_{a}$. Consider now theories of the form:
    $\tilde{\mathcal{L}}_{F_{ab}\mathcal{F}_{ab}}=\hat{\mathcal{L}}(\mathfrak{s},\mathfrak{p})+\mathcal{L}(s,p,\lambda\mathfrak{p})\,$, (60)
    where $\mathfrak{s}=-\frac{1}{4}\mathcal{F}_{ab}\mathcal{F}^{ab}$, $\hat{\mathcal{L}}$ is an arbitrary function of the indicated variables and where $\mathcal{L}(s,p,\psi)$ satisfies (6). The equation of motion for $\mathcal{F}_{ab}$ is given by:
    $\partial_{a}\left(\frac{\partial\hat{\mathcal{L}}}{\partial\mathfrak{s}}\mathcal{F}^{ab}\right)+\partial_{a}\left(\frac{\partial\hat{\mathcal{L}}}{\partial\mathfrak{p}}+\lambda\,\frac{\partial\mathcal{L}}{\partial\psi}\right)\star\mathcal{F}^{ab}=0\,.$ (61)
    Again, we directly see that (61) is exactly invariant under electromagnetic duality rotations. One could additionally wonder whether there exist theories $\hat{\mathcal{L}}(\mathfrak{s},\mathfrak{p})$ for which the complete theory $\tilde{\mathcal{L}}_{F_{ab}\mathcal{F}_{ab}}$ is also invariant under duality rotations associated with the extra gauge field $\mathcal{A}_{a}$. This lies beyond the scope of the present manuscript and is planned to be studied elsewhere.
  3. 3.
    $\psi=\lambda\chi_{\rm GB}$, where $\lambda$ is a dimensionless coupling and $\chi_{\rm GB}=R_{abcd}R^{abcd}-4R_{ab}R^{ab}+R^{2}$ stands for the Gauss-Bonnet density. Take the following theory:
    $\tilde{\mathcal{L}}_{F_{ab},g_{ab}}=\frac{R}{16\pi G_{N}}+\mathcal{L}(s,p,\lambda\chi_{\rm GB})\,$, (62)
    where $\mathcal{L}(s,p,\psi)$ fulfills (6). Define the following tensor [42] $P_{ab}{}^{cd}$:
    $P_{ab}{}^{cd}=\frac{\partial\tilde{\mathcal{L}}_{F_{ef},g_{ef}}}{\partial R_{cd}{}^{ab}}=\frac{1}{16\pi G_{N}}\delta_{[a}^{c}\delta_{b]}^{d}+2\lambda\frac{\partial\mathcal{L}}{\partial\psi}\Sigma_{ab}{}^{cd}\,,\quad\Sigma_{ab}{}^{cd}=R_{ab}{}^{cd}-4R_{[a}^{[c}\delta_{b]}^{d]}+R\delta_{[a}^{c}\delta_{b]}^{d}\,.$ (63)
    The gravitational equations of motion for the theory $\tilde{\mathcal{L}}_{F_{ab},g_{ab}}$ will be given by [43]:
    $P_{(a}{}^{cde}R_{b)cde}-\frac{1}{2}g_{ab}\tilde{\mathcal{L}}_{F_{ef},g_{ef}}+2\nabla_{c}\nabla_{d}P_{(a}{}^{c}{}_{b)}{}^{d}+\frac{\partial\mathcal{L}}{\partial F^{(a}{}_{c}}F_{b)c}=0\,.$ (64)
    We compute:
    $P_{(a}{}^{cde}R_{b)cde}=\frac{R_{ab}}{16\pi G_{N}}+\frac{\lambda}{2}\frac{\partial\mathcal{L}}{\partial\psi}\chi_{\rm GB}g_{ab}\,,\quad 2\nabla_{c}\nabla_{d}P_{(a}{}^{c}{}_{b)}{}^{d}=4\lambda\nabla_{c}\nabla_{d}\left(\frac{\partial\mathcal{L}}{\partial\psi}\right)\Sigma_{a}{}^{c}{}_{b}{}^{d}\,$$\quad\,$. (65)
    As a result, (64) simplifies down to:
    $\frac{G_{ab}}{16\pi G_{N}}+\frac{1}{2}g_{ab}\left[\lambda\frac{\partial\mathcal{L}}{\partial\psi}\chi_{\rm GB}-\mathcal{L}\right]+4\lambda\nabla_{c}\nabla_{d}\left(\frac{\partial\mathcal{L}}{\partial\psi}\right)\Sigma_{a}{}^{c}{}_{b}{}^{d}+\frac{\partial\mathcal{L}}{\partial F^{(a}{}_{c}}F_{b)c}=0\,.$ (66)
    Now, let us explore how this equation transforms under an infinitesimal duality rotation of angle $\alpha$. Direct computation shows:
    $\delta_{\alpha}\left(\lambda\frac{\partial\mathcal{L}}{\partial\psi}\chi_{\rm GB}-\mathcal{L}\right)=\alpha\left(-\lambda\chi_{\rm GB}+2p\right)\,,\quad\delta_{\alpha}\left(\frac{\partial\mathcal{L}}{\partial F^{(a}{}_{c}}F_{b)c}\right)=\alpha g_{ab}\left(-p+\frac{\lambda\chi_{\rm GB}}{2}\right)\,.$ (67)
    As a consequence, we clearly observe that (66) is invariant under duality rotations. In any event, we also note that the (66) will include higher derivatives of the metric and the gauge field.
Appendix C Solving (6) and (35) perturbatively